# Machine learning surrogates for efficient hydrologic modeling: Insights from stochastic simulations of managed aquifer recharge

### v1

@@ 3.1.1 Impact of training set size on accuracy
While the PredRNN++ exhibits the lowest E2E error, it also demonstrates the second lowest sensitivity to training set size. Figure 7 reveals that the ML models, in general, exhibit a power-law relationship between training set size and MAPE, which is typical for most deep learning architectures (Hestness et al., 2017). This power-law relationship demonstrates diminishing returns when increasing training set size beyond a certain point. To explore these diminishing returns more closely, we plot the derivative of the best-fit curve for this power-law relationship ($\frac{d\text{MAPE}}{dm}(m)$) in Figure 7. The $\frac{d\text{MAPE}}{dm}(m)$ values for the PredRNN++ remain consistently close to 0 and the model is able to achieve $\sim$5MAPE with as few as 25 training examples (Figure 7). In fact, almost all models exhibit nearly consistent MAPE values after 50–100 training examples. While the models presented in Figure 5 were trained on 245 examples (with a 217-28 training-validation split), these results suggests that future practitioners could achieve similar performance by training on fewer examples. However, the relationship between training set size and MAPE is highly dependant on the specific problem of interest, so future practitioners should use caution when extending these results to other sites or other scenarios.
@@ 3.1.2 Impact of normalization on accuracy
While normalizing input data improves model accuracy, loss normalization increases training stability. Comparing MAPE on the validation set across all training epochs, we see that using a normalized loss generally reduces the number of sharp spikes in MAPE, indicating that a normalized loss allows for more stable descent and reliable training convergence. One reason loss normalization may have such a large impact in this case is our use of small batch sizes. To limit memory requirements, we train each ML model using a batch size of 1. However, at small batch sizes, the gradient is calculated over fewer examples (or in this case, a single example), which leads to noisy loss curves and increases the likelihood that gradient descent can get stuck in a local minimum. Loss normalization helps offset these drawbacks by ensuring that the gradients are of consistent magnitude across the training set. This effect is most noticeable in models that use either no data normalization or Z-score data normalization, suggesting that loss normalization most benefits models with inadequate data normalization.
@@ 3.2 Computational efficiency
While each ML surrogate model achieves $<$$10\%$ MAPE, the models vary widely in the amount of time and memory they require to produce simulations. One attribute they share, however, is that all ML surrogate models offer significant runtime advantages over the process-based hydrologic model (Figure 9). Figure 9 reveals positive slopes across all lines. Recall that the slope for these plots are calculated as the difference in time taken by ParFlow-CLM and the ML model in generating a single Stage 1 simulation (Equation 6). All lines appear to exhibit identical slopes because each ML surrogate outputs simulations orders of magnitude faster than the process-based hydrologic model ($t_{\text{PF}}=$ 15,017 CPU-seconds/simulation), so the difference in $t_{\text{PF}}-t_{\text{ML}}$ across the ML models is miniscule. Nonetheless, positive slopes for all lines signify that all seven ML models produce simulations at a faster rate than ParFlow-CLM. The ViT4d technically exhibits the steepest slope, producing simulations the fastest ($t_{\text{ML}}=0.115$ GPU-seconds/simulation), but, due to the large training set size required for the ViT4d, time savings only appear for the ViT4d after 246 simulations. Closer examination of the $x$-intercepts reveals that 6 out of 7 of the ML models achieve a runtime advantage over ParFlow-CLM when the desired simulation output count is as low as 17. We provide a detailed overview of all time measurements in Table 2, including distinct values of $\mathcal{D}_{\text{train}}$ (the minimum size training set required to achieve $<10\%$ MAPE, which strongly impacts the computational efficiency of each model.
Figure 9: The time saved, in processor-hours, when performing a given number of simulations using an ML model as compared to using ParFlow-CLM. $T_{\text{saved}}$ is calculated following Equation 6 and includes both training time and evaluation time. Note that the U-FNO4d line is almost entirely overlapped by the PredRNN++ line.
@@
    The minimum number of training examples required to achieve $<$$10\%$ MAPE. We show the training-validation split in parentheses.
@@ 3.3 Model interpretability
Another crucial factor to consider when selecting an ML architecture within a hybrid modeling framework is model interpretability. Error recognition is more straightforward with an interpretable surrogate model and can enhance trust among practitioners. The PredRNN++ architecture, for example, contains $\mathcal{C}$ and $\mathcal{M}$ memory cells (Figure S2) that are intended to serve distinct roles, with the former facilitating the transfer of temporal information and the latter facilitating the transfer of spatial information (Wang et al., 2018). However, in practice, deciphering the values stored in these cells and validating their role in each model has proven challenging (Figure 10). Each memory cell is the same size as the model domain ($120\times 80\times 25$), with two memory cells for each of the four causal LSTM layers in the PredRNN++. Given that these memory cells change between each time step and between each evaluation example, manually inspecting and interpreting $\mathcal{C}$ and $\mathcal{M}$ quickly becomes cumbersome. In other words, the hidden states of the PredRNN++ model and, consequently, the internal mechanisms of the model, prove challenging to interpret.
@@
In addition to the PredRNN++, other ML architectures have been shown to be interpretable when applied in different domains and may prove useful in hydrologic surrogate modeling. For example, previous studies have indicated that hierarchical convolutional layers, as employed in CNN3d, can learn hierarchical image information (Zhou et al., 2014; Yang et al., 2019). Furthermore, research on transformers like those employed by the ViT3d has demonstrated that attention mechanisms within these models may also learn hierarchical information (Jain and Wallace, 2019; Jawahar et al., 2019; Wiegreffe and Pinter, 2019). When combined with conventional ML interpretability methods such as saliency maps, deconvolutional networks, guided backpropagation, and feature occlusion, each of these mechanisms could serve as a check to ensure that a surrogate model is behaving as expected. Future work is needed to explore the interpretablility of each of these architectures when applied to hydrologic simulations.
@@ 3.4 Limitations
The full results that display ML model performance on all stages in Figure 5 reveal a similar challenge in capturing temporal changes in pressure head. The performance in Stage 1 for all models consistently lags behind that of Stage 2 and Stage 3. Because these later stages simulate shorter time periods (one year in Stage 2 and two years in Stage 3, as opposed to 440 years in Stage 1), pressure head changes over the model stage are less pronounced and their temporal relationships are thereby easier to learn. Thus, while this case study achieves low MAPE, future applications should carefully adapt the implementation strategies in this study (for example, the choice of loss function, data normalization technique and input features) to their problem of interest, especially when simulating model domains characterized by abrupt variations in pressure head across either time or space.
@@ 4 Conclusion
Machine learning surrogate models offer an efficient approach for generating large sets of hydrologic simulations. All seven of the ML surrogate models produce transient pressure fields that are comparable (below 10% mean absolute percentage error, MAPE) to those produced by the process-based hydrologic model. These transient pressure fields can then be used in downstream tasks, such as quantifying changes in groundwater storage. The improved predictive recurrent neural network (PredRNN++) exhibits the highest accuracy when simulating all three model stages (0.36% MAPE), though other architectures also perform well (0.52–2.4% MAPE). Interestingly, he non-sequential models, which are deeper, have more parameters and compute pressure at all time steps simultaneously, do not exhibit improved accuracy over the one-step models, which compute pressure for each time step sequentially and with no memory between steps.
This hybrid modeling framework also yields substantial runtime benefits (0.115–2.38 processor-seconds/simulation, as opposed to 3,860–69,700 processor-seconds/simulation with a process-based model for Stage 1 simulations) with errors consistently below 10%. Even when accounting for the time it takes to train each model, 6 out of 7 ML surrogate models exhibit runtime advantages when generating as few as 17 hydrologic simulations. Computational savings increase when performing larger numbers of simulations, with 159$\times$ speed-up when generating 500 simulations with the U-FNO4d, for example.
Additionally, we identify key practices for efficiently training accurate hydrologic surrogate models, showing that min-max data normalization and loss normalization increase surrogate model accuracy and training stability. Temporal and spatial downsampling, meanwhile, improve computational efficiency by decreasing the memory required to train each model.
@@ Acknowledgements
This supplementary file contains additional details on the methods used to develop each ML surrogate model, including layer-by-layer information for each surrogate model (§S1.1), diagrams of recent model architecture innovations (§S1.2), and details on the autoencoder used to compress 3D pressure fields (§S1.3). We also present additional results referenced in the main text, including complete MAPE results across all models and stages (§S2.1) and a discussion of any error introduced by the autoencoder (§S2.2).
S1 Materials and methods
@@ 7. Linear $(-1,120,80,25,1)$ 33
Figure S2: PredRNN++ (Wang et al., 2018) architecture diagram.

### v2

@@ 3.1.1 Impact of training set size on accuracy
While the PredRNN++ exhibits the lowest E2E error, it also demonstrates the second lowest sensitivity to training set size. Figure 7 reveals that the ML models, in general, exhibit a power-law relationship between training set size and MAPE, which is typical for most deep learning architectures (Hestness et al., 2017). This power-law relationship demonstrates diminishing returns when increasing training set size beyond a certain point. To explore these diminishing returns more closely, we plot the derivative of the best-fit curve for this power-law relationship ($\frac{d\text{MAPE}}{dm}(m)$) in Figure 7. The $\frac{d\text{MAPE}}{dm}(m)$ values for the PredRNN++ remain consistently close to 0, indicating a low sensitivity to changes in training set size, and the model is able to achieve $\sim$5MAPE with as few as 25 training examples (Figure 7). In fact, almost all models exhibit nearly consistent MAPE values after 50–100 training examples. While the models presented in Figure 5 were trained on 245 examples (with a 217-28 training-validation split), these results suggests that future practitioners could achieve similar performance by training on fewer examples. In §3.2, we even find that some hydrologic surrogate models such as the U-FNO4d and PredRNN++ can achieve $<$10% MAPE with as low as 3 training examples. However, the relationship between training set size and MAPE is highly dependent on the specific problem of interest, so future practitioners should use caution when extending these results to other sites or other scenarios.
@@ 3.1.2 Impact of normalization on accuracy
While normalizing input data improves model accuracy, loss normalization increases training stability. Comparing MAPE on the validation set across all training epochs, we see that using a normalized loss generally reduces the number of sharp spikes in MAPE, indicating that a normalized loss allows for more stable descent and reliable training convergence. One reason loss normalization may have such a large impact in this case is our use of small batch sizes. To prevent memory overflow in the PredRNN++, we train with a batch size of 1; we use a batch size of 1 for all other models to maintain consistent training configurations. However, at small batch sizes, the gradient is calculated over fewer examples (or in this case, a single example), which leads to noisy loss curves and increases the likelihood that gradient descent can get stuck in a local minimum. Loss normalization helps offset these drawbacks by ensuring that the gradients are of consistent magnitude across the training set. This effect is most noticeable in models that use either no data normalization or Z-score data normalization, suggesting that loss normalization most benefits models with inadequate data normalization.
@@ 3.2 Computational efficiency
While each ML surrogate model achieves $<$$10\%$ MAPE, the models vary widely in the amount of time and memory they require to produce simulations. One attribute they share, however, is that all ML surrogate models offer significant runtime advantages over the process-based hydrologic model (Figure 9). Figure 9 reveals positive slopes across all lines. Recall that the slope for these plots are calculated as the difference in time taken by ParFlow-CLM and the ML model in generating a single Stage 1 simulation (Equation 6). All lines appear to exhibit identical slopes because each ML surrogate outputs simulations orders of magnitude faster than the process-based hydrologic model ($t_{\text{PF}}=$ 15,017 CPU-seconds/simulation), so the difference in $t_{\text{PF}}-t_{\text{ML}}$ across the ML models is miniscule. Nonetheless, positive slopes for all lines signify that all seven ML models produce simulations at a faster rate than ParFlow-CLM. The ViT4d technically exhibits the steepest slope, producing simulations the fastest ($t_{\text{ML}}=0.115$ GPU-seconds/simulation), but, due to the large training set size required for the ViT4d to achieve $<$10% MAPE, time savings only appear for the ViT4d after 246 simulations. Closer examination of the $x$-intercepts reveals that 6 out of 7 of the ML models achieve a runtime advantage over ParFlow-CLM when the desired simulation output count is as low as 17. 2 out of the 7, the U-FNO4d and PredRNN++, achieve a runtime advantage on and after the fourth simulation. We provide a detailed overview of all time measurements in Table 2, including distinct values of $\mathcal{D}_{\text{train}}$ (the minimum size training set required to achieve $<10\%$ MAPE), which strongly impacts the computational efficiency of each model.
Figure 9: The time saved, in processor-hours, when performing a given number of simulations using an ML model as compared to using ParFlow-CLM. (b) is a local enlargement of (a), with the enlarged region labeled with “Inset” in (a). $T_{\text{saved}}$ is calculated following Equation 6 and includes both training time and evaluation time. Note that the U-FNO4d line is almost entirely overlapped by the PredRNN++ line.
@@
    The minimum number of training examples required to achieve $<$$10\%$ test MAPE. We show the training-validation split in parentheses.
@@ 3.3 Model interpretability
Another crucial factor to consider when selecting an ML architecture within a hybrid modeling framework is model interpretability. Error recognition is more straightforward with an interpretable surrogate model and can enhance trust among practitioners, ensuring that a surrogate model is behaving as expected. The PredRNN++ architecture, for example, contains $\mathcal{C}$ and $\mathcal{M}$ memory cells (Figure S2) that are intended to serve distinct roles, with the former facilitating the transfer of temporal information and the latter facilitating the transfer of spatial information (Wang et al., 2018). However, in practice, deciphering the values stored in these cells and validating their role in each model has proven challenging (Figure 10). Each memory cell is the same size as the model domain ($120\times 80\times 25$), with two memory cells for each of the four causal LSTM layers in the PredRNN++. Given that these memory cells change between each time step and between each evaluation example, manually inspecting and interpreting $\mathcal{C}$ and $\mathcal{M}$ quickly becomes cumbersome. In other words, the hidden states of the PredRNN++ model and, consequently, the internal mechanisms of the model, prove challenging to interpret.
@@
Other ML architectures have been shown to be more interpretable than the PredRNN++ when applied in different domains and may prove useful in hydrologic surrogate modeling, such as the CNN3d, which employs convolutional layers, capable of learning hierarchical image information (Zhou et al., 2014; Yang et al., 2019), or the ViT3d, whose attention mechanisms can also learn hierarchical information (Jain and Wallace, 2019; Jawahar et al., 2019; Wiegreffe and Pinter, 2019). Future work is needed to explore the interpretablility of each of these architectures when applied to hydrologic simulations.
@@ 3.4 Limitations
The full results that display ML model performance on all stages in Figure 5 reveal a similar challenge in capturing temporal changes in pressure head. The performance in Stage 1 for all models consistently lags behind that of Stage 2 and Stage 3. Because these later stages simulate shorter time periods (one year in Stage 2 and two years in Stage 3, as opposed to 440 years in Stage 1), pressure head changes over the model stage are less pronounced and their temporal relationships are thereby easier to learn. Thus, while this case study achieves low MAPE, future applications should carefully adapt the implementation strategies in this study (for example, the choice of loss function, data normalization technique, and input features) to their problem of interest, especially when simulating model domains characterized by abrupt variations in pressure head across either time or space.
Finally, while we have examined the impact of modeling choices such as loss function and data normalization across several surrogate architectures, our analysis has been limited to a single case study. Future research should apply these techniques to a broader range of hydrologic scenarios to further validate and generalize these findings.
@@ 4 Conclusion
Machine learning surrogate models offer an efficient approach for generating large sets of hydrologic simulations. All seven of the ML surrogate models produce transient pressure fields that are comparable (below 10% mean absolute percentage error, MAPE) to those produced by the process-based hydrologic model. These transient pressure fields can then be used in downstream tasks, such as quantifying changes in groundwater storage. The improved predictive recurrent neural network (PredRNN++) exhibits the highest accuracy when simulating all three model stages (0.36% MAPE), though other architectures also perform well (0.52–2.4% MAPE). To achieve $<$10% MAPE for this case study, some architectures such as the U-FNO4d and PredRNN++ require $<$10 training examples.
This hybrid modeling framework yields substantial runtime benefits, though the threshold at which it becomes more efficient than a process-based model varies by architecture. Once trained, each architecture requires 0.115–2.38 processor-seconds to generate a single Stage 1 simulation, as opposed to the process-based model, which requires 3,860–69,700 processor-seconds/simulation. When accounting for the time it takes to train each model, 6 out of 7 ML surrogate models (all except for the four-dimensional vision transformer) exhibit runtime advantages when generating as few as 17 hydrologic simulations. Because they require so few training examples, the U-FNO4d and PredRNN++ achieve runtime advantages on and after the fourth simulation. These findings indicate that, when generating more than 17 hydrologic simulations for this case study, a surrogate model is more computationally efficient than a process-based model, even when including the time required to generate a training set and train the surrogate. In contrast, when performing fewer than four simulations for this case study, a process-based model alone is more efficient.
Additionally, we identify the impacts of key modeling choices on surrogate accuracy, showing that min-max data normalization can improve surrogate model accuracy by up to a factor of 10 when compared to other treatments such as Z-score and no normalization. Selecting a normalized loss function also improves training stability. Downsampling input features using an autoencoder, meanwhile, reduces memory requirements by training with tensors 4% of their original size.
@@ Acknowledgements
All process-based hydrologic simulations were performed with ParFlow-CLM v3.10.0 (https://doi.org/10.5281/zenodo.6413322; Smith et al., 2022). Output files from process-based hydrologic simulations are available to the public via the Stanford Digital Repository (https://doi.org/10.25740/hj302gv2126; Perzan et al., 2023b). All code used to implement, train and evaluate each machine learning surrogate architecture is available to the public in HydroShare (https://doi.org/10.4211/hs.f0a31fbc3de148a98deb36795b4fac53; Dai et al., 2025).
@@
This appendix contains additional details on the methods used to develop each ML surrogate model, including layer-by-layer information for each surrogate model (§S1.1), diagrams of recent model architecture innovations (§S1.2), and details on the autoencoder used to compress 3D pressure fields (§S1.3). We also present additional results referenced in the main text, including complete MAPE results across all models and stages (§S2.1) and a discussion of any error introduced by the autoencoder (§S2.2).
Appendix S1 Materials and methods
@@ 7. Linear $(-1,120,80,25,1)$ 33
Figure S2: PredRNN++ (Wang et al., 2018) architecture diagram. The memory cells in the blue highlighted region are saved for the next time step. As the forward pass for the next time step is being computed, the model will utilize the saved memory cells from the previous forward pass. Note that the temporal $\mathcal{C}$ memory is not passed to deeper layers within the model; it is retained within the same causal LSTM layer across each time step.